# Cross-Layer Vision Smoothing: Enhancing Visual Understanding via Sustained Focus on Key Objects in Large Vision-Language Models

### v1

@@ B.1 Results on MME
We evaluated the performance of our method on a wider range of tasks using the MME benchmark, as shown in Table 5. Our method achieves the highest overall score among all baselines, but exhibits relatively weak performance on the position perception task. We argue that the position task relies heavily on spatial information in the image, and our approach—by prolonging LVLMs’ focus on key objects—does not enhance their understanding of spatial relationships. In contrast, our method significantly improves LVLMs’ comprehension of key objects in the color, count, and OCR tasks.
B.2 Results on POPE
The effectiveness of CLVS stems from the LVLM’s inherent attention to key objects. However, when no key objects are present in the image—such as in the negative samples of object hallucination tests—CLVS becomes ineffective. The results on POPE clearly illustrate this issue: our method significantly improves recall, i.e., accuracy on positive samples, but performs poorly on negative samples, resulting in unstable improvement.
Figure 8: Effects of $\gamma$ in LLaVA-Next and Qwen2.5-VL.
B.3 $\gamma$~ for LLaVA-Next and Qwen2.5-VL
$\gamma$ is a hyperparameter that acts as a soft window size, limiting how far the model can look back into previous layers. We search for the optimal value of $\gamma$, and the experimental results are shown in Figure 8. Based on the results, we set $\gamma=0.7$ for LLaVA-Next and Qwen2.5-VL.

### v2

@@ B Results on MME
We evaluated the performance of our method on a wider range of tasks using the MME benchmark [MME], as shown in Table 5. Our method achieves the highest overall score among all baselines, but exhibits relatively weak performance on the position perception task. We argue that the position task relies heavily on spatial information in the image, and our approach—by prolonging LVLMs’ focus on key objects—does not enhance their understanding of spatial relationships. In contrast, our method significantly improves LVLMs’ comprehension of key objects in the color, count, and OCR tasks.
C Effectiveness on Other MLLMs
We validate the effectiveness of our method on the Qwen series of models. Since Qwen2.5-VL [qwen25-vl] employs 3D positional indices, we disable the Unified Position module in our experiments with Qwen2.5-VL. The results in Table 6 show that CLVS still achieves significant improvements on the Qwen series models, further verifying the positive impact of sustained visual attention in vision understanding.